# Masked Auto-Regressive Variational Acceleration: Fast Inference Makes Practical Reinforcement Learning

### v1

@@ Abstract
Masked auto-regressive diffusion models (MAR) benefit from the expressive modeling ability of diffusion models and the flexibility of masked auto-regressive ordering. However, vanilla MAR suffers from slow inference due to its hierarchical inference mechanism: an outer AR unmasking loop and an inner diffusion denoising chain. Such decoupled structure not only harm the generation efficiency but also hinder the practical use of MAR for reinforcement learning (RL), an increasingly critical paradigm for generative model post-training. To address this fundamental issue, we introduce MARVAL (Masked Auto-regressive Variational Acceleration), a distillation-based framework that compresses the diffusion chain into a single AR generation step while preserving the flexible auto-regressive unmasking order. Such a distillation with MARVAL not only yields substantial inference acceleration but, crucially, makes RL post-training with verifiable rewards practical, resulting in scalable yet human-preferred fast generative models. Our contributions are twofold: (1) a novel score-based variational objective for distilling masked auto-regressive diffusion models into a single generation step without sacrificing sample quality; and (2) an efficient RL framework for masked auto-regressive models via MARVAL-RL. On ImageNet 256×256, MARVAL-Huge achieves an FID of 2.00 with more than 30 times speedup compared with MAR-diffusion, and MARVAL-RL yields consistent improvements in CLIP and image-reward scores on ImageNet datasets with entity names. In conclusion, MARVAL demonstrates the first practical path to distillation and RL of masked auto-regressive diffusion models, enabling fast sampling and better preference alignments.
@@ 1 Introduction
Auto-regressive (AR) [van2016pixel, reed2017few, van2016conditional, reed2017parallel, tian2024visual] and diffusion-based [dhariwal2021diffusion, song2020score] generative models exhibit complementary strengths and weaknesses. Diffusion models excel at high-fidelity synthesis and mode coverage, but require long denoising trajectories with hundreds of iterative updates [ho2020denoising, salimans2016improved, song2020denoising], making inference and downstream Reinforcement Learning (RL) post-training computationally expensive. In contrast, AR models generate sequentially in discrete space, offering natural support for semantic controllability and compatibility with RL and language-based alignment frameworks [esser2021taming, ramesh2021zero, luong2015effective], but they often struggle to match the perceptual quality of state-of-the-art diffusion models at high resolutions. Motivated by their complementary properties, recent works have combined the two paradigms—forming AR+Diffusion [hoogeboom2021autoregressive, gu2022vector, li2024autoregressive] architectures that use AR priors to capture global structure and diffusion decoders to restore fine-grained details. The Masked Auto-regressive (MAR) model [li2024autoregressive] represents a prominent example of this hybrid design.
However, MAR models face two fundamental limitations. First, their nested architecture—where each AR iteration contains a full diffusion unmasking process—results in extremely slow inference. Second, this inefficiency significantly hinders practices to perform RL-based post-training, which is essential for scalable generative models to be aligned with human preferences[christiano2017deep, ouyang2022training, stiennon2020learning]. The above two issues encourage us to answer a fundamental research question: Can we improve the efficiency of MAR-like models while making RL post-training practical and efficient?
@@
Empirically, MARVAL-H achieves competitive ImageNet [deng2009imagenet] $256\times 256$ generation quality (FID = 2.00), which slightly outperforms MAR-H (FID = 2.06) while being 20× faster. When post-trained with RL using human preference rewards, MARVAL-RL further improves CLIP and Image Reward scores, demonstrating better alignment with human preferences. As illustrated in Fig. 1, MARVAL-RL delivers visually sharper and semantically richer generations than the MAR baseline while maintaining real-time inference efficiency ($\approx 0.6$ s per image). Together, these results establish MARVAL as a practical and scalable path toward efficient, RL-ready image generative modeling with a masked auto-regressive generation order.
@@ 2.1 Diffusion Models
Diffusion models [ho2020denoising, song2020score, song2020denoising] try to train a neural network to model data distribution by gradually corrupting samples with a Gaussian forward diffusion process. From a continuous-time perspective, Song et al[song2020score] generalized diffusion models using stochastic differential equations (SDEs). The forward perturbation process is an Itô SDE:
@@
which explicitly involves the score function $\nabla_{\boldsymbol{x}_{t}}\log p_{t}(\boldsymbol{x}_{t})$. Since this quantity is unknown, it is approximated by a neural network $\boldsymbol{S}_{\theta}(\boldsymbol{x}_{t},t)$ by minimizing a weighted denoising score matching loss [song2019generative]:
@@
where the weighting function $\lambda(t)$ balances different noise levels. Once trained, the score network can be directly integrated into Eq. 2 to simulate reverse SDEs [song2020score, karras2022elucidating, xue2023sa] or ODEs [song2020denoising, lu2022dpm], thereby generating high-quality samples.
@@ 2.2 Masked Auto-regressive Model
The most common practice of traditional diffusion models tries to model the entire image simultaneously in some latent space [kingma2013auto]. However, such a mechanism brings difficulties in modeling complex correlations at potentially high dimensions. Li et al. [li2024autoregressive] proposed Masked Auto-Regressive Models (MAR), which brings flexible random generation orders by introducing masked auto-regressive ordering. MAR first transforms an image latent representation into n tokens $(x_{1},...,x_{n})$. Then it randomly masks some tokens in the sequence, and uses the unmasked tokens as conditions to train a lightweight diffusion head to model the denoising process of unmasked tokens.
@@ 2.3 One-step Diffusion and Preference Alignment
A major challenge of diffusion models is their slow multi-step sampling process, which requires hundreds to thousands of network evaluations. To address this, various distillation techniques have been proposed to accelerate sampling while preserving high-quality generation [salimans2022progressive, song2023consistency, luo2024diff, yin2023one, zhou2024score, wang2025uni, kim2023consistency, songimproved, zhou2024long, heek2024multistep, xie2024distillation, salimans2024multistep, geng2024consistency, meng2022distillation, sauer2023adversarial, xu2024ufogen, zheng2024trajectory, li2024reward, berthelot2023tract, luo2023comprehensive]
Score Implicit Matching (SIM) [luo2024one] is a recent and powerful approach that enables one-step distillation of a pre-trained diffusion model without requiring access to real data during distillation. Unlike traditional knowledge distillation, which requires a dataset of teacher outputs, SIM operates by directly matching the score functions of the teacher and the student generator. Let the teacher’s score function at noise level $t$ be $s_{p}(\boldsymbol{x},t)$, and the generator-induced score be $s_{q}(\boldsymbol{x},t)$. SIM minimizes a time-integrated score divergence between the two:
@@
where $d(\cdot,\cdot)$ is a distance function (e.g., squared distance) and $\pi_{t}$ is a reference sampling distribution. A key insight of SIM is that while the generator’s score $s_{q}$ is intractable, the gradient of this objective with respect to the generator’s parameters $\theta$ can be efficiently computed. This is achieved by training an auxiliary network $\boldsymbol{S}_{\phi}(\boldsymbol{x_{t}},t)$ to approximate $s_{q}$, and then minimizing a tractable surrogate objective derived from the score-divergence gradient theorem  [luo2024one].
@@
With the improved techniques of one-step diffusion, several works have explored practices of human preference alignment of diffusion models[dai2023emu, podell2023sdxl, prabhudesai2023aligning, clark2023directly, lee2023aligning, fan2024reinforcement, black2023training, wallace2024diffusion, yang2024using, hong2024margin]. Particularly, some recent works [luo2025reward, luo2024diffpp, luo2024david, luo2024diffstar] have advanced the human preference alignment for one-step diffusion models, with rigorous results. However, considering the introduction of random masked auto-regressive ordering, how to properly conduct preference alignment for one-step masked auto-regressive models remains unexplored.
@@ 3.1 Guided Score Implicit Matching (GSIM)
Therefore, we introduce Guided Score Implicit Matching (GSIM). Inspired by the research of Uni-instruct [wang2025uni], the KL divergence is the integral of the Fisher divergence along a diffusion process under mild regularity conditions. Consequently, Eq. 7 can be expressed in continuous time as the integral over $t$ of the score discrepancy $D_{\text{score}}$ between the student’s time-dependent score $\nabla_{\mathbf{x}_{t}}\log q_{\theta,t}(\mathbf{x}_{t}\mid c)$ and the pretrained diffusion model’s score $s_{p_{t}}(\mathbf{x}_{t},c)$:
@@
Since MAR employs classifier-free guidance (CFG) [ho2022classifier] during inference, the teacher score $s_{p_{t}}(\boldsymbol{x}_{t},c)$ in Eq. 8 must correspond to the CFG-guided diffusion process. Accordingly, we define the guided teacher score as:
@@
where $s_{q_{\text{sg}[\theta],t}}(\boldsymbol{x}_{t},c)$ can be approximated by the auxiliary network $\boldsymbol{S}_{\phi}(\boldsymbol{x}_{t},t,c)$. Although theory suggests using the squared $\ell_{2}$ distance $\|\cdot\|_{2}^{2}$ for $d(\cdot,\cdot)$ in Eq. 8 is necessary to recover the KL objective exactly, in practice this choice can cause training instability. For improved training robustness and faster convergence, we adopt the Pseudo-Huber distance as the loss function in our implementation, following [luo2024one]. Concretely, in Eq. 11 we set
@@ 3.2 Reinforcement Learning for Post-Hoc Refinement
By back propagating the gradient from this reward signal, we fine-tune $\theta$ to produce outputs that are not only consistent with the auto-regressive prior but also explicitly optimized for higher quality and better text-image alignment. For our specific implementation, we utilize PickScore [kirstain2023pick] as the reward model $R(\cdot)$, as it is a powerful function known to correlate well with human aesthetic judgments and preferences.
@@ 4 Experiments
In this section, we evaluate the proposed distillation and reinforcement learning (RL) scheme for our MARVAL (Masked Auto-regressive Variational Acceleration). We conduct three sets of experiments to investigate (1) the influence of classifier-free guidance (CFG) scales on the distilled model, (2) the acceleration achieved over MAR under varying numbers of autoregressive iterations and diffusion steps, (3) the overall effectiveness and significance of RL refinement across models of different scales.
@@ 4.1 Experimental Setup
Model Settings
Following MAR [li2024autoregressive], our diffusion process adopts a cosine noise schedule with 1000 training steps. The denoising network predicts the noise vector $\varepsilon$ through a lightweight MLP. We evaluate three model scales the same as MAR, denoted as Base, Large, and Huge, which differ in the depth and width of the denoising MLP used in the diffusion component. The details of different models are illustrated in Table 1. In the distillation stage, we train each model on 8 NVIDIA A800 GPUs for 30 epochs, which takes approximately 3 days to complete. In the RL refinement stage, we further refine the distilled model on the same hardware configuration for 5 epochs, requiring about 2 days of training.
Model Type MLP Depth MLP Width #Params
Base 6 1024 208M
Large 8 1280 479M
Huge 12 1536 943M
Table 1: Different scales of MAR model with diffusion loss. Depth and width refer to the number of residual MLP blocks and hidden channel dimensions respectively.
Evaluation Metrics.
We evaluate two aspects of performance:
  • Text–image alignment: measured by CLIP Score [hessel2021clipscore] and ImageReward Score [xu2023imagereward], using prompts in the form of “a high-quality and harmonious picture of + class name”.
  • Image quality: measured by Fréchet Inception Distance(FID) [heusel2017gans] and Inception Score (IS) [salimans2016improved].
In order to ensure the robustness and accurate results, all our metrics are tested on 50K generated samples uniformly distributed in 1000 classes corresponding to ImageNet dataset.
@@ 4.4 The Significance of Reinforcement Learning
To assess the impact of reinforcement learning (RL) fine-tuning, we perform an ablation study comparing models before and after RL optimization. Starting from the distilled model, RL fine-tuning is applied to better align image generation with perceptual and semantic objectives. RL markedly improves visual quality and perceptual realism, as reflected in the significant improvent of reward-based metrics in Table 4.
@@ 5 Conclusion
Limitations and Future Directions
Our work enables efficient generation and practical RL post-training for MAR models, but its RL stage still requires multi-step AR iterations together with a large reward model. This overhead imposes substantial memory demands, which consequently restricts the feasible batch size in the RL stage. This effect becomes increasingly pronounced as the model scales up, leading to less efficient RL refinement. A practical mitigation is to shrink the active training graph, such as using lighter reward models or reducing AR iterations appropriately to relieve memory pressure and improve efficiency.
A natural extension of our distillation + RL framework is applying it to other AR + diffusion architectures beyond the class-conditional image generation of MAR. In particular, the same principles could benefit large text-to-image systems and multi-modal AR models, enabling faster sampling and more efficient preference optimization. Exploring these directions may provide a broader validation of the generality and cross-modal applicability of our approach.
[…]
@@ 1 Theory Part
In this section we present a detailed proof of the Gradient Equivalent Theorem. The proof of the Gradient Equivalent Theorem is based on the so-called Score-projection identity which was first found in  [vincent2011connection] to bridge denoising score matching and denoising auto-encoders. Later the identity is applied for deriving distillation methods based on Fisher divergences  [zhou2024score]. We appreciate the efforts of Zhou et al.  [zhou2024score] and re-write the score-projection identity here without proof. Readers can check Zhou et al.  [zhou2024score] for a complete proof of score-projection identity.

### v2

@@ Abstract
Masked auto-regressive diffusion models (MAR) benefit from the expressive modeling ability of diffusion models and the flexibility of masked auto-regressive ordering. However, vanilla MAR suffers from slow inference due to its hierarchical inference mechanism: an outer AR unmasking loop and an inner diffusion denoising chain. Such decoupled structure not only harm the generation efficiency but also hinder the practical use of MAR for reinforcement learning (RL), an increasingly critical paradigm for generative model post-training. To address this fundamental issue, we introduce MARVAL (Masked Auto-regressive Variational Acceleration), a distillation-based framework that compresses the diffusion chain into a single AR generation step while preserving the flexible auto-regressive unmasking order. Such a distillation with MARVAL not only yields substantial inference acceleration but, crucially, makes RL post-training with verifiable rewards practical, resulting in scalable yet human-preferred fast generative models. Our contributions are twofold: (1) a novel score-based variational objective for distilling masked auto-regressive diffusion models into a single generation step without sacrificing sample quality; and (2) an efficient RL framework for masked auto-regressive models via MARVAL-RL. On ImageNet 256×256, MARVAL-Huge achieves an FID of 2.00 with more than 30 times speedup compared with MAR-diffusion, and MARVAL-RL yields consistent improvements in CLIP and image-reward scores on ImageNet datasets with entity names. In conclusion, MARVAL demonstrates the first practical path to distillation and RL of masked auto-regressive diffusion models, enabling fast sampling and better preference alignments. Project page: https://ai4imaginglab.github.io/MARVAL-RL-Page/
@@ 1 Introduction
Auto-regressive (AR) [57, 41, 56, 42, 55] and diffusion-based [8, 52] generative models exhibit complementary strengths and weaknesses. Diffusion models excel at high-fidelity synthesis and mode coverage, but require long denoising trajectories with hundreds of iterative updates [15, 44, 48], making inference and downstream Reinforcement Learning (RL) post-training computationally expensive. In contrast, AR models generate sequentially in discrete space, offering natural support for semantic controllability and compatibility with RL and language-based alignment frameworks [9, 40, 34], but they often struggle to match the perceptual quality of state-of-the-art diffusion models at high resolutions. Motivated by their complementary properties, recent works have combined the two paradigms—forming AR+Diffusion [18, 13, 25] architectures that use AR priors to capture global structure and diffusion decoders to restore fine-grained details. The Masked Auto-regressive (MAR) model [25] represents a prominent example of this hybrid design.
However, MAR models face two fundamental limitations. First, their nested architecture—where each AR iteration contains a full diffusion unmasking process—results in extremely slow inference. Second, this inefficiency significantly hinders practices to perform RL-based post-training, which is essential for scalable generative models to be aligned with human preferences[4, 36, 53]. The above two issues encourage us to answer a fundamental research question: Can we improve the efficiency of MAR-like models while making RL post-training practical and efficient?
@@
Empirically, MARVAL-H achieves competitive ImageNet [7] $256\times 256$ generation quality (FID = 2.00), which slightly outperforms MAR-H (FID = 2.06) while being 20× faster. When post-trained with RL using human preference rewards, MARVAL-RL further improves CLIP and Image Reward scores, demonstrating better alignment with human preferences. As illustrated in Fig. 1, MARVAL-RL delivers visually sharper and semantically richer generations than the MAR baseline while maintaining real-time inference efficiency ($\approx 0.6$ s per image). Together, these results establish MARVAL as a practical and scalable path toward efficient, RL-ready image generative modeling with a masked auto-regressive generation order.
@@ 2.1 Diffusion Models
Diffusion models [15, 52, 48] try to train a neural network to model data distribution by gradually corrupting samples with a Gaussian forward diffusion process. From a continuous-time perspective, Song et al. [52] generalized diffusion models using stochastic differential equations (SDEs). The forward perturbation process is an Itô SDE:
@@
which explicitly involves the score function $\nabla_{\boldsymbol{x}_{t}}\log p_{t}(\boldsymbol{x}_{t})$. Since this quantity is unknown, it is approximated by a neural network $\boldsymbol{S}_{\theta}(\boldsymbol{x}_{t},t)$ by minimizing a weighted denoising score matching loss [51]:
@@
where the weighting function $\lambda(t)$ balances different noise levels. Once trained, the score network can be directly integrated into Eq. 2 to simulate reverse SDEs [52, 19, 63] or ODEs [48, 26], thereby generating high-quality samples.
@@ 2.2 Masked Auto-regressive Model
The most common practice of traditional diffusion models tries to model the entire image simultaneously in some latent space [21]. However, such a mechanism brings difficulties in modeling complex correlations at potentially high dimensions. Li et al. [25] proposed Masked Auto-Regressive Models (MAR), which brings flexible random generation orders by introducing masked auto-regressive ordering. MAR first transforms an image latent representation into n tokens $(x_{1},...,x_{n})$. Then it randomly masks some tokens in the sequence, and uses the unmasked tokens as conditions to train a lightweight diffusion head to model the denoising process of unmasked tokens.
@@ 2.3 One-step Diffusion and Preference Alignment
A major challenge of diffusion models is their slow multi-step sampling process, which requires hundreds to thousands of network evaluations. To address this, various distillation techniques have been proposed to accelerate sampling while preserving high-quality generation [45, 49, 27, 65, 68, 60, 20, 50, 67, 14, 61, 46, 12, 35, 47, 62, 66, 24, 2, 31]
Score Implicit Matching (SIM) [28] is a recent and powerful approach that enables one-step distillation of a pre-trained diffusion model without requiring access to real data during distillation. Unlike traditional knowledge distillation, which requires a dataset of teacher outputs, SIM operates by directly matching the score functions of the teacher and the student generator. Let the teacher’s score function at noise level $t$ be $s_{p}(\boldsymbol{x},t)$, and the generator-induced score be $s_{q}(\boldsymbol{x},t)$. SIM minimizes a time-integrated score divergence between the two:
@@
where $d(\cdot,\cdot)$ is a distance function (e.g., squared distance) and $\pi_{t}$ is a reference sampling distribution. A key insight of SIM is that while the generator’s score $s_{q}$ is intractable, the gradient of this objective with respect to the generator’s parameters $\theta$ can be efficiently computed. This is achieved by training an auxiliary network $\boldsymbol{S}_{\phi}(\boldsymbol{x_{t}},t)$ to approximate $s_{q}$, and then minimizing a tractable surrogate objective derived from the score-divergence gradient theorem  [28].
@@
With the improved techniques of one-step diffusion, several works have explored practices of human preference alignment of diffusion models[6, 38, 39, 5, 23, 11, 3, 59, 64, 17]. Particularly, some recent works [33, 32, 29, 30] have advanced the human preference alignment for one-step diffusion models, with rigorous results. However, considering the introduction of random masked auto-regressive ordering, how to properly conduct preference alignment for one-step masked auto-regressive models remains unexplored.
@@ 3.1 Guided Score Implicit Matching (GSIM)
Therefore, we introduce Guided Score Implicit Matching (GSIM). Inspired by the research of Uni-instruct [60], the KL divergence is the integral of the Fisher divergence along a diffusion process under mild regularity conditions. Consequently, Eq. 7 can be expressed in continuous time as the integral over $t$ of the score discrepancy $D_{\text{score}}$ between the student’s time-dependent score $\nabla_{\mathbf{x}_{t}}\log q_{\theta,t}(\mathbf{x}_{t}\mid c)$ and the pretrained diffusion model’s score $s_{p_{t}}(\mathbf{x}_{t},c)$:
@@
Since MAR employs classifier-free guidance (CFG) [16] during inference, the teacher score $s_{p_{t}}(\boldsymbol{x}_{t},c)$ in Eq. 8 must correspond to the CFG-guided diffusion process. Accordingly, we define the guided teacher score as:
@@
where $s_{q_{\text{sg}[\theta],t}}(\boldsymbol{x}_{t},c)$ can be approximated by the auxiliary network $\boldsymbol{S}_{\phi}(\boldsymbol{x}_{t},t,c)$. Although theory suggests using the squared $\ell_{2}$ distance $\|\cdot\|_{2}^{2}$ for $d(\cdot,\cdot)$ in Eq. 8 is necessary to recover the KL objective exactly, in practice this choice can cause training instability. For improved training robustness and faster convergence, we adopt the Pseudo-Huber distance as the loss function in our implementation, following [28]. Concretely, in Eq. 11 we set
@@ 3.2 Reinforcement Learning for Post-Hoc Refinement
By back propagating the gradient from this reward signal, we fine-tune $\theta$ to produce outputs that are not only consistent with the auto-regressive prior but also explicitly optimized for higher quality and better text-image alignment. For our specific implementation, we utilize PickScore [22] as the reward model $R(\cdot)$, as it is a powerful function known to correlate well with human aesthetic judgments.
@@ 4 Experiments
In this section, we evaluate the proposed distillation and reinforcement learning (RL) scheme for our MARVAL (Masked Auto-regressive Variational Acceleration). We conduct four sets of experiments to investigate (1) the influence of classifier-free guidance (CFG) scales on the distilled model, (2) the acceleration achieved over MAR under varying numbers of autoregressive iterations and diffusion steps, (3) the overall effectiveness and significance of RL refinement across models of different scales, and (4) the generalization capability of our framework to text-to-image generation tasks.
@@ 4.1 Experimental Setup
Following MAR [25], our diffusion process adopts a cosine noise schedule with 1000 training steps. The denoising network predicts the noise vector $\varepsilon$ through a lightweight MLP. We evaluate three model scales identical to MAR: Base (6 blocks, 1024 width, 208M params), Large (8 blocks, 1280 width, 479M params), and Huge (12 blocks, 1536 width, 943M params). In the distillation stage, we train each model on 8 NVIDIA A800 GPUs for 30 epochs, which takes approximately 3 days to complete. In the RL refinement stage, we further refine the distilled model on the same hardware configuration for 5 epochs, requiring about 2 days of training.
For robust evaluation, class-conditional metrics are evaluated on 50K samples uniformly distributed across 1,000 ImageNet classes, while text-to-image metrics use 50K prompts randomly sampled from the JourneyDB [54] test set.
@@ 4.4 The Significance of Reinforcement Learning
To assess the impact of reinforcement learning (RL) fine-tuning, we perform an ablation study comparing the distilled models before and after RL optimization. RL fine-tuning is applied to better align image generation with perceptual and semantic objectives. RL markedly improves visual quality and perceptual realism, as reflected in the significant improvent of reward-based metrics in Table 3.
@@
4.5 Generalization to Text-to-Image Generation
To demonstrate our framework’s versatility beyond class-conditional generation, we apply our MARVAL-RL framework to the Text-to-Image (T2I) DC-AR model [10].
T2I tasks require high-fidelity synthesis and strict prompt alignment. Evaluated on 50,000 samples (Table 4), our 1-step+RL strategy compresses the original 20-step DDIM sampling into a single step. The compression does not degrade performance; instead, RL fine-tuning significantly boosts the ImageReward score from 0.5876 to 0.6903, highlighting RL’s value in complex alignment tasks.
Qualitatively (Figure 5), the 1-step+RL DC-AR model maintains excellent visual and semantic fidelity. It accurately captures fine-grained details (e.g., facial structure, hair texture) and adheres strictly to the prompt, proving our framework successfully preserves generative quality even under extreme acceleration. Additional high-quality generation results are provided in the supplementary material.
Model (Text-to-Image) Strategy ImageReward $\uparrow$
DC-AR-512 20-step DDIM 0.5876
DC-AR (distilled+RL) 1-step + RL 0.6903
Table 4: ImageReward on 50K T2I samples. Our strategy compresses inference while significantly boosting alignment.
Figure 5: T2I qualitative comparison. Left: Original 20-step DC-AR. Right: Our 1-step+RL DC-AR. The RL-finetuned one-step model exhibits enhanced perceptual quality and visual harmony.
@@ 1 Theory Part
In this section we present a detailed proof of the Gradient Equivalent Theorem. The proof of the Gradient Equivalent Theorem is based on the so-called Score-projection identity which was first found in  [58] to bridge denoising score matching and denoising auto-encoders. Later the identity is applied for deriving distillation methods based on Fisher divergences  [68]. We appreciate the efforts of Zhou et al.  [68] and re-write the score-projection identity here without proof. Readers can check Zhou et al.  [68] for a complete proof of score-projection identity.